Problem 3 A city is evaluating a proposal to erect and operate a parking facility in its downtown area. Three designs on avaiAnswer is in bold. Need to know how to work the problem. Thanks!
Answer #1

B/C Ratio = PV of annual fee / (Cost of site + Cost of building + PV of annual maintenance cost)

Therefore,

B/C Ratio for Design 1 = [830 x P/A(10%, 30)] / [240 + 2,200 + 410 x P/A(10%, 30)]

= (830 x 9.4269) / [2,440 + (410 x 9.4269)]

= 7,824 / (2,440 + 3,865)

= 7,824 / 6,305

= 1.24

B/C Ratio for Design 2 = [600 x P/A(10%, 30)] / [180 + 700 + 360 x P/A(10%, 30)]

= (600 x 9.4269) / [880 + (360 x 9.4269)]

= 5,656 / (880 + 3,394)

= 5,656 / 4,274

= 1.32

B/C Ratio for Design 3 = [620 x P/A(10%, 30)] / [200 + 1,400 + 310 x P/A(10%, 30)]

= (620 x 9.4269) / [1,600 + (310 x 9.4269)]

= 5,845 / (1,600 + 2,922)

= 5,845 / 4,522

= 1.29

Since Design 2 has highest B/C Ratio, this is the best alternative.

NOTE: "Choose 1" is not the answer that's written - it means question is asking you to choose 1 option out of the given 3.